Find: 3 whole ½ ÷ 4 Find: 3 whole ½ ÷ 4 Find: 3 whole ½ ÷ 4

 
Find:
3 whole ½ ÷ 4
Find:
3 whole ½ ÷ 4
Find:
3 whole ½ ÷ 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First covert the mixed fraction into improper fraction. We have,
= 3 whole 1⁄2 = 7/2
Then,
= (7/2) × reciprocal of 4
= (7/2) × (1/4)
= (7 × 1) / (2 × 4)
= 7/8

Thanks